OpenStudy (anonymous):

Computer Craft pays its sales associates a 15% commission on every sale. Barry sold a laptop computer for $1,900 and a printer for $300. Question:How much money did Barry earn on the sale of the items? Choices:A. $285 B. $330 C. $345 D. $550

OpenStudy (anonymous):

15% of $1,900 = 0.15 * 1900 = $285 15% of $300 = 0.15 * 300 = $45 285 + 45 = $330